# Advances in Robust Federated Learning: Heterogeneity Considerations

### v1

@@ III-B2 Model Compression
Model compression refers to compressing the size of a model and is commonly used to solve model, device and communication heterogeneity issues. The commonly used compression methods in FL include sparsification, quantization and factorization. Sparsification usually means identify and retain the meaningful enough parameters, while setting other parameters to some constants. In this way, the parameter matrices are transformed to be sparse, and only those meaningful enough parameters are transmitted. Different from sparsification, quantization keeps the parameters dense and focuses on reducing the bit-width occupation of each parameter. A representative quantization method is SignSGD [91], which only retains the sign bit of each parameter. Factorization decompose a large matrix into the product of multiple low-rank matrices or vectors, and thus compress the size of the whole model with limited precision degradation. To compress both upstream and downstream communications, Sattler et al. [92] point out that sparsification is more robust to Non-IID data and thus propose Sparse Ternary Compression (STC), which further quantizes the remaining top-$k$ elements to the mean population magnitude. Similar to the goal of [92], Amiri et al. [93] propose Lossy Federated Learning (LFL). The server broadcast the quantized version the global model parameters to the clients, and the uploading parameters are also quantized. Shah et al. [94] propose to compress both downstream and upstream communications, but it goes further. To recover the loss of sparsity after aggregation, each client can perform sparse reconstruction for the averaged model, and it can also directly begin local training with the received sparse model, which can help in inducing sparsity in the client model. Recently, Tang et al. [95] point out that SignSGD is vulnerable to Non-IID data, and current corresponding methods cannot support multiple local SGD updates like FedAvg. Thus, they propose to inject some random noise before the sign operation to control the bias brought by it, and further propose the first sign-based FedAvg algorithm $z$-SignFedAvg.
To combat the challenge of model and device heterogeneity, Yao et al. [96] propose FedHM, where the server compresses the large global model via low-rank factorization for each client to satisfy its resource availability, while preserving the model’s performance. Moreover, the compressed model can be transformed back to full-rank model for aggregation. Jeong et al. [97] propose Factorized-FL, which aims to factorize model parameters into basis vectors. Such factorized parameters space can make aggregation operation more robust to different labels and data, and separate client-specific and client-general knowledge. Recently, to perform personalized FL under scenario with data heterogeneity and device heterogeneity, Chen et al. [98] propose a novel personalized FL framework Fed-QSSL, where clients can perform low-bitwidth quantization training locally to satisfy its device constraint, and tackle data heterogeneity with distributed self-supervised learning.
@@ III-B3 Knowledge Distillation
Knowledge distillation [99] is originally proposed for transferring knowledge from a larger model to a smaller model, and thus reducing the number of parameters. Generally speaking, This process requires the output of the smaller model to approach the larger one. Therefore, knowledge distillation can also be used for knowledge transferring between participants in FL. To alleviate the impact of feature skew and model heterogeneity, Huang et al. [100] propose FCCL, where each client constructs a cross-correlation matrix according to other clients’ soft labels on public dataset. Besides, FCCL also utilize dual-domain knowledge distillation to alleviate catastrophic forgetting caused by domain shift. Zhu [101] et al. propose FedGEN, where the server trains a light-weighted representation generator, and broadcast it to clients. While local training, the clients utilize the generated representations for knowledge distillation, hence get rid of reliance on public datasets. For scenarios with model heterogeneity and quality skew, Fang et al. [38] proposes RHFL, where clients combine cross entropy and reverse cross entropy to alleviate the impact of label noise, and perform knowledge distillation on a public dataset to reduce the knowledge gap between clients. Besides, RHFL also introduces client confidence re-weighting to reduce the negative impact caused by other clients’ noise. Recently, to combat severe label noise because of highly contaminated clients, Lu et al. [23] propose FedNed, which identifies noisy clients and requires them to train models with noisy labels and pseudo-labels simultaneously. Then the models trained with noisy labels act as bad teachers for negative distillation, while the ones trained with pseudo-labels are used for aggregation. To learn personalized models under data heterogeneity, Chen et al. [102] propose a method, which captures similarity between generic model and personalized models by spectral distillation. Moreover, it also establishes a co-distillation framework to bridge the knowledge of the generic model and personalized models.
@@ III-B4 Federated Optimization
Federated multi-task learning. The core idea of multi-task learning is to train a model simultaneously to perform multiple related tasks, thereby improving the model’s generalization ability. By regarding each client’s local training as a separate task, multi-task learning techniques can help clients to learn personalized model with high generalization ability. Smith et al. [9] extend MTL method CoCoA from traditional distributed learning to federated environment, and propose MOCHA, which considers each client’s local training as a task and train a personalized model for each client. Corinzia et al. [103] propose to construct a Bayesian network between the server and the clients, and utilize variational inference for its optimization. Comparing to MOCHA, it requires no additional regularization term in local training objective and thus improve the efficiency of local training and ensuring network-agnostic FL. Besides, it also extends FMTL to scenarios with non-convex objectives. To learn a personalized model for devices that later join FL, Li et al. [104] propose OFMTL, which helps the new devices to learn with the assistance of weight matrix and precision matrix constructed by old devices, requiring no presence of earlier devices. Dinh et al. [105] propose FedU, which introduce Laplacian regularization to determine the correlation between clients in FMTL. Besides, They also propose a decentralized version dFedU. Marfoq et al. [106] consider the data distribution of each client as a mixture of several latent distributions. Therefore, they propose FedEM, where clients can collaboratively train several generic component models, and each client only needs to learn how to mix these components to build a personalized model.
@@
Meta-learning based FL. Meta-learning, which is also known as ”learning to learn”, is usually used to design and train algorithms to quickly adapt to new tasks, typically when it is very little data. The core goal of meta-learning is to improve the flexibility and adaptability of learning algorithms, enabling them to learn from previous experiences and better handle new and unknown tasks. Chen et al. [107] combine FL with meta-learning and then propose FedMeta, where all clients collaboratively train a initialization model. In FedMeta, each client upload its test loss on its query set which is used to update the initialization model. Each client can learn a personalized model from the initialization model by just a few local training, thus FedMeta can also reduce the training burden of the clients. Recently, Jeon et al. [108] propose a Bayesian meta-learning based approach called MetaVD, where the server establishes a shared hypernetwork for predicting client-specific dropout rates, which are utilized for few-shot local adaptation on clients. Leo et al. [109] propose FedL2P, which introduces meta-nets to inductively map each client’s Batch Normalization statistics to client-specific personalization hyperparameters. Scott et al. [110] propose PeFLL. The server is equipped with a hypernetwork, which takes the descriptor of each client as input and personalized model parameters as output. The descriptor of each client is obtained with an embedding network, taking some data points as input. Thereby PeFLL overcomes the problems of high latency and client-side computation cost.
@@
Regularization based FL. Regularization usually means adding a regularization term in the local training objectives and thus achieving some specific goals, such as tackling data heterogeneity. Li et al. [43] propose FedProx, which adds a proximal term in local training objective to reduce the distance between the local model and the global model, and thus handles data heterogeneity. Besides, They also propose the concept of in-exact solution, enabling the clients to upload incompletely solved parameters and thus alleviate the problem of device heterogeneity. Kim et al. [111] propose FedMLB, which separate models into several blocks, and each client form a main pathway and several hybrid pathways with the blocks. Regularization terms based on knowledge distillation are formed with the output of the main pathway and the hybrid pathways. Shoham et al. [112] point out that due to data heterogeneity, the problem of forgetting in Sequential Lifelong Learning may also occur in FL. Therefore, they propose FedCurv, which introduces Elastic Weight Consolidation (EWC) to regularize model parameters during local training, and thus avoid forgetting the learned information from other clients. Similarly, Yao et al. [113] propose FedCL, which utilize a proxy dataset to construct importance weight matrix of all local models to perform EWC regularization, and thus alleviate weight divergence. Dinh et al. [114] propose pFedME, which introduce Moreau Envelope for regularization. It adds an L2 regularization term to the local objective to transform the entire FL into a two-level objective. The inner optimization obtains optimal personalized models while the outer one obtains the global model. Karimireddy et al. [115] propose SCAFFOLD, utilizing control variables that reflect the update direction of the global model and the local models respectively, and then correct update direction of local model with a regularization term formed by their differences. Hanzely et al. [116] propose a new formulation of FL, which no longer learn a single global model, but a mixture of the global model and the pure personalized models. To perform such FL formulation, this works proposes a novel randomized gradient-based method L2GD, for minimizing the loss function composed of average loss and penalty term. Li et al. [117] propose a personalized FL method Ditto, where each client perform local training on the global model and then on its personalized model, and applies regularization to the personalized model’s training to bring it closer to the global model. Besides, They also theoretically and empirically prove that Ditto has good robustness and fairness guarantee. Recently, Zhou et al. [118] propose FedLNL, which designs a novel method based on Bayesian Inference for updating local Noise Transition Matrix (NTM) and thus avoid too large gradients, and then apply a diversity product regularizer in local training to utilize this method. Zhi et al. [119] suggest that each client can guide its local training referring to other clients’ parameters, and thus propose a regularizer based on a relation cube to conduct layer-wise parameters coaching. Li et al. [120] empirically observe that current FL algorithms are highly susceptible to the degree of weight decay, and propose FedNAR. FedNAR is an algorithmic plug-in which can help controlling the magnitude of weight decay.
@@ III-C Architecture-Level Methods
To address the federated heterogeneity challenge, making adjustments in the training structure is an emerging improvement. In this section, we discuss some heterogeneous methods that make adjustments on the FL architecture. Specifically, we classify these structure-level approaches into client selection, clustered federated learning, hierarchical federated learning, and decentralized federated learning.
@@ III-C1 Client selection
Client selection refers to the process of selecting a subset of available clients to participate in model updates during the training process. This selection can directly impact the efficiency and effectiveness of the training. Effective client selection can address issues of data heterogeneity and constraints related to communication and computation [121, 122].
To alleviate the pressure of heterogeneity, Nishio et al. [53] proposed that FedCS uses a greedy algorithm to solve the client selection problem with resource constraints, which allows the server to aggregate as many client updates as possible and accelerate the performance improvement of ML models. Yoshida et al. [123] adopted a heuristic algorithm and proposed a hybrid learning mechanism called Hybrid-FL, trying to select the best set of clients to use their own data to train the model. Wang et al. [7] introduced an experience-driven control framework, Favor, which intelligently selects client devices to participate in each round of FL to offset the bias introduced by non-IID data and accelerate the convergence speed. Xu et al. [124] proposes a stochastic optimization problem that jointly connects client selection and bandwidth allocation under long-term client constraints. lai2021oort et al. [125] proposed Oort to improve the performance of joint training and testing by guiding participant selection. Oort uses existing data to improve model accuracy and speed up training. Wu et al. [126] proposed a probabilistic node selection framework, FedPNS, which dynamically changes the probability of each node being selected based on the output of the optimal aggregation. Cho et al. [127] presents a convergence analysis of FL of biased customer selection via efficient selection aggregation and quantifies how bias affects the convergence speed. Li et al. [128] proposed PyramidFL to speed up FL training while achieving higher final model performance. At the heart of PyramidFL is fine-grained client selection, not only focusing on the differences in client selection between those selected and unselected participants, but also fully exploiting data and system heterogeneity within selected clients. Ribero et al. [129] proposes a federated averaging algorithm assisted by an adaptive sampling technique, F3AST, an unbiased algorithm that dynamically learns an availability-dependent client selection strategy that asymptotically minimizes client sampling variance to global model convergence The impact improves the performance of joint learning.
@@ III-C2 Clustered Federated Learning
Clustering FL is used to deal with situations where the data distribution is highly heterogeneous. This approach trains the model in a more homogeneous group by clustering clients based on the similarity of their data.
CFL (Cluster Federated Learning) [130] is a method of recursively clustering clients by exploiting the cosine similarity between client gradient updates during each round of FL training. This clustering allows for a more targeted and efficient distribution of training tasks among clients with similar characteristics. IFCA (Intra-cluster Federated Average) [131] is another clustered FL method that utilizes the loss of models within different clusters on the client’s local dataset to cluster clients. In each round, IFCA randomly selects a portion of clients within each cluster for training and updates the intra-cluster models by aggregating the local models within the cluster. PACFL (Personalized and Clustered Federated Learning) [132] clusters clients by uploading data information from each client to the server and taking measurements of the principal angles between client data subspaces before the FL process begins. During each round of joint learning, PACFL updates each intra-cluster model with a weighted aggregation of the local models of the intra-cluster clients, enabling personalized learning while still taking advantage of clustering. FedSoft [133] is a cluster FL framework designed to train both local personalized models and high-quality cluster models. In order to limit the workload of the clients, FedSoft employs proximal updating, which allows only a subset of clients to complete an optimization task in each round of communication. ACFL (Auction-based Cluster Federated Learning) [134] introduces a mean-shift clustering algorithm that intelligently groups clients based on their local data distributions. This clustering approach allows FL systems to identify groups of clients with similar data characteristics, enabling more targeted and efficient model training.
III-C3 Hierarchical Federation Learning

### v2

@@ III-B2 Model Compression
Model compression methods are aiming at compressing the size of a model and is commonly used to solve model, device and communication heterogeneity issues, including sparsification, quantization and factorization. Sparsification usually means identify and retain the meaningful enough parameters, while setting other parameters to some constants. In this way, the parameter matrices are transformed to be sparse, and only those meaningful enough parameters are transmitted. Different from sparsification, quantization keeps the parameters dense and focuses on reducing the bit-width occupation of each parameter. A representative quantization method is SignSGD [114], which only retains the sign bit of each parameter. Factorization decompose a large matrix into the product of multiple low-rank matrices or vectors, and thus compress the size of the whole model with limited precision degradation. To compress both upstream and downstream communications, Sattler et al. [58] point out that sparsification is more robust to Non-IID data and thus propose Sparse Ternary Compression (STC), which further quantizes the remaining top-$k$ elements to the mean population magnitude and only transfer the remaining elements. Similar to the goal of [58], Amiri et al. [115] propose Lossy FL (LFL). The server broadcast the quantized version the global model parameters to the clients, and the uploading parameters are also quantized. Shah et al. [59] propose to compress both downstream and upstream communications, but it goes further. To recover the loss of sparsity after aggregation, each client can perform sparse reconstruction for the averaged model, and it can also directly begin local training with the received sparse model. Recently, Tang et al. [116] point out that SignSGD is vulnerable to Non-IID data, and current corresponding methods cannot support multiple local SGD updates like FedAvg. Thus, they propose to inject some random noise before the sign operation to control the bias brought by it, and further propose the first sign-based FedAvg algorithm $z$-SignFedAvg.
To combat the challenge of model and device heterogeneity, Yao et al. [63] propose FedHM, where the server compresses the large global model via low-rank factorization for each client to satisfy its resource availability. The compressed model can be transformed back to full-rank model for aggregation. Jeong et al. [43] propose Factorized-FL, which factorize model parameters into basis vectors. It can make aggregation operation more robust to different labels and data, helping address the issue of data heterogeneity. Recently, Chen et al. [64] propose a personalized FL framework Fed-QSSL, where clients can perform low-bitwidth quantization training locally to satisfy its device constraint, and tackle data heterogeneity with distributed self-supervised learning.
@@ III-B3 Knowledge Distillation
Knowledge distillation [117] is originally proposed for transferring knowledge from a larger model to a smaller model, and thus reducing the number of parameters. Generally speaking, This process requires the output of the smaller model to approach the larger one. Therefore, KD can also be used for knowledge transferring between participants in FL.
Knowledge distillation is mostly adopted for handling model heterogeneity. Li et al. [9] firstly consider model heterogeneity in FL and propose FedMD. FedMD introduces a public dataset, on which the clients upload their private models’ soft labels to the server for aggregation as consensus. The clients utilize the consensus as the output of the teacher model and transfer it to their own private model by KD on public dataset, and then perform a few epochs of training on their own private dataset. Therefore, the clients can safely share their knowledge without exposing model structure and transmitting parameters. Cronus proposed by chang et al. [118] is similar to FedMD, but it takes into account the possibility of attacks during aggregating the knowledge and improves the process of aggregation to be robust. Li et al. [119] propose FedH2L, where the clients can learn knowledge from others by accessing their public datasets and perform KD, and correct the gradients to reduce the conflicts of gradients between private and public dataset caused by data heterogeneity. Cheng et al. [120] consider how to train a larger global model in the case of heterogeneous client models and propose FedGEMS, which selective upload client outputs on public dataset, and establish a selective knowledge fusion mechanism on the server for effective knowledge transfer. To alleviate the impact of feature skew and model heterogeneity, Huang et al. [41] propose FCCL, where each client constructs a cross-correlation matrix according to other clients’ soft labels on public dataset, and utilize dual-domain KD to alleviate catastrophic forgetting caused by domain shift. For scenarios with model heterogeneity and quality skew, Fang et al. [45] proposes RHFL, where clients combine cross entropy and reverse cross entropy to alleviate the impact of label noise, and perform KD on a public dataset to reduce the knowledge gap between clients.
The above works all introduce public datasets, but in reality, such datasets are not available in most cases. Therefore, some other works mainly focus on utilizing KD without relying on public datasets. Shen et al. [51] propose FML, which shares a small meme model between the server and all clients. Each client performs mutual learning between its private model and the meme model, and then upload the parameters of the meme model for aggregation. Therefore, knowledge can be shared between the clients through the meme model. Wu et al. [48] propose FedKD, which also utilizes a global model for KD on logit and representation space, and the communication cost of global model is reduced by performing SVD-decomposition on its parameters. He et al. [65] propose FedGKT for training large CNNs when the performance of edge devices is limited. The clients upload computed features and soft labels as well as real labels to the server. The server utilize them to conduct KD, and send its soft labels over the features back to the clients for local KD. FedDKC proposed by Wu et al. [121] is similar to FedGKT, but they propose two mechanisms for client knowledge congruence to mitigate the impact of knowledge differences between clients. Zhang et al. [122] propose FedGD, which introduces a distributed trained generator fixed on the server to generate intermediate features. The clients upload classifier parameters and soft labels for generator training, and perform KD on the generated intermediate features. Zhang et al. [123] also propose a generator-based method called FedDTG. Unlike FedGD, in FedDTG, the generator directly generates samples in raw data space and is not fixed on the server, but is trained by adversarial training with local classifiers and discriminators on the client side.
@@
Apart from model heterogeneity, knowledge distillation can also be utilized for handling other types of heterogeneity. To alleviate the impact of data heterogeneity, Yao et al. [124] propose FedGKD, where the server stores global model parameters aggregated in recent multiple historical rounds. During local training, the knowledge from the historical global models is transferred to the client models through KD. Zhang et al. [29] propose FedFTG, where the server maintains a generator trained with clieng models for generating pseudo data, and the server can perform KD based on the pseudo data to transfer knowledge from the client models to the aggregated global model. Zhu et al. [125] propose FedGEN, where the server trains a light-weighted generator for representation generation. While local training, the clients utilize the generated representations for KD. Recently, to combat severe label noise caused by highly contaminated clients, Lu et al. [24] propose FedNed, which identifies noisy clients and requires them to train models with noisy labels and pseudo-labels simultaneously. Then the models trained with noisy labels act as bad teachers for negative distillation, while the ones trained with pseudo-labels are used for aggregation. To learn personalized models under data heterogeneity, Chen et al. [30] propose a method to capture similarity between generic model and personalized models by spectral distillation. Moreover, it establishes a co-distillation framework to bridge the knowledge of the generic model and personalized models.
@@ III-B4 Federated Optimization
Federated multi-task learning. The core idea of multi-task learning is to train a model simultaneously to perform multiple related tasks, thereby improving the model’s generalization ability. By regarding each client’s local training as a separate task, multi-task learning techniques can help clients to learn personalized model with high generalization ability. Smith et al. [10] extend MTL method CoCoA from traditional distributed learning to federated environment, and propose MOCHA, which considers each client’s local training as a task. MOCHA learns a separate model for each client instead of a global model, which is robust to data heterogeneity. Corinzia et al. [126] propose to construct a Bayesian network between the server and the clients, and utilize variational inference for its optimization. Comparing to MOCHA, it requires no additional regularization term in local training objective and thus ensuring network-agnostic FL. Besides, it also extends FMTL to scenarios with non-convex objectives. To learn a personalized model for devices that later join FL, Li et al. [127] propose OFMTL, which helps the new devices to learn with the assistance of weight matrix and precision matrix constructed by old devices, requiring no presence of earlier devices. Dinh et al. [128] propose FedU, which introduce Laplacian regularization to determine the correlation between clients in FMTL. Besides, They also propose a decentralized version dFedU. Marfoq et al. [129] consider the data distribution of each client as a mixture of several latent distributions. Therefore, they propose FedEM, where clients can collaboratively train several generic component models, and each client only needs to learn how to mix these components to build a personalized model.
@@
Meta-learning based FL. Meta-learning, which is also known as ”learning to learn”, is usually used to design and train algorithms to quickly adapt to new tasks, typically when it is very little data. The core goal of meta-learning is to improve the flexibility and adaptability of learning algorithms, enabling them to learn from previous experiences and better handle new and unknown tasks. Chen et al. [31] combine FL with meta-learning and then propose FedMeta, where all clients collaboratively train a initialization model. Each client upload its test loss on its query set which is used to update the initialization model. Each client can learn a personalized model from the initialization model by just a few local training, thus FedMeta can also reduce the training burden of the clients and reduce the performance degradation of the final personalized models due to the scenarios with data heterogeneity. Recently, Jeon et al. [130] propose a Bayesian meta-learning based approach called MetaVD, where the server establishes a shared hypernetwork for predicting client-specific dropout rates, which are utilized for few-shot local adaptation on clients. Leo et al. [131] propose FedL2P, which introduces meta-nets to inductively map each client’s Batch Normalization statistics to client-specific personalization hyperparameters. Scott et al. [132] propose PeFLL. The server is equipped with a hypernetwork, which takes the descriptor of each client as input and personalized model parameters as output. The descriptor of each client is obtained with an embedding network, taking some data points as input. Thereby PeFLL overcomes the problems of high latency and client-side computation cost. However, when the required model’s size is significantly large, such hypernetwork could be hard to establish because there too many model parameters to output.
@@
Regularization based FL. Regularization usually means adding a regularization term in the local training objectives and thus achieving some specific goals, such as tackling data heterogeneity. Li et al. [32] propose FedProx, which adds a proximal term in local training objective to reduce the distance between the local model and the global model, and thus handles data heterogeneity. Besides, They also propose in-exact solution, enabling the clients to upload incompletely optimized parameters and thus alleviate the problem of device heterogeneity. Kim et al. [38] propose FedMLB, which separate models into several blocks, and each client form a main pathway and several hybrid pathways with the blocks. Regularization terms based on KD are formed with the output of the main pathway and the hybrid pathways. By performing KD on hybrid pathways between the global model and the local model, FedMLB can bridge the knowledge gap caused by data heterogeneity, especially distribution skew and label skew. Shoham et al. [133] point out that due to data heterogeneity, the problem of forgetting may also occur in FL. Therefore, they propose FedCurv, which introduces Elastic Weight Consolidation (EWC) to regularize model parameters, and thus avoid forgetting the learned information from other clients. Similarly, Yao et al. [134] propose FedCL, which utilize a proxy dataset to construct importance weight matrix of all local models to perform EWC regularization, and thus alleviate weight divergence. Dinh et al. [135] propose pFedME, which introduce Moreau Envelope for regularization. It adds an L2 regularization term to the local objective to transform the entire FL into a two-level objective. The inner optimization obtains optimal personalized models while the outer one obtains the global model. Karimireddy et al. [33] propose SCAFFOLD, utilizing control variables that reflect the update direction of the global model and the local models respectively, and then correct update direction of local model with a regularization term formed by their differences, which can help alleviate the performance degradation caused by data heterogeneity. Li et al. [136] propose a personalized FL method Ditto, where each client perform local training on the global model and then on its personalized model, and applies regularization to the personalized model’s training to bring it closer to the global model. Besides, They also theoretically and empirically prove that Ditto has good robustness and fairness guarantee. Recently, Zhou et al. [137] propose FedLNL, which designs a novel method based on Bayesian Inference for updating local Noise Transition Matrix (NTM) and thus avoid too large gradients, and then apply a diversity product regularizer in local training to utilize this method. Zhi et al. [34] suggest that each client can guide its local training referring to other clients’ parameters, and thus propose a regularizer based on a relation cube to conduct layer-wise parameters coaching. Li et al. [138] empirically observe that current FL algorithms are highly susceptible to the degree of weight decay, and propose FedNAR. FedNAR is an algorithmic plug-in which can help controlling the magnitude of weight decay.
@@ III-C Architecture-Level Methods
To address the federated heterogeneity challenge, making adjustments in the training structure is an emerging improvement. In this section, we discuss some heterogeneous methods that make adjustments on the FL architecture. Specifically, we classify these structure-level approaches into client selection, clustered FL, hierarchical FL, and decentralized FL.
@@ III-C1 Client selection
Client selection refers to the process of selecting a subset of available clients to participate in model updates during the training process. This selection can directly impact the efficiency and effectiveness of the training. Effective client selection can address issues of data heterogeneity and constraints related to communication and computation [139].
To alleviate the pressure of heterogeneity, Nishio et al. [96] proposed that FedCS uses a greedy algorithm to solve the client selection problem with resource constraints, which allows the server to aggregate as many client updates as possible and accelerate the performance improvement of ML models. Yoshida et al. [140] adopted a heuristic algorithm and proposed a hybrid learning mechanism called Hybrid-FL, trying to select the best set of clients to use their own data to train the model. Wang et al. [7] introduced an experience-driven control framework, Favor, which intelligently selects client devices to participate in each round of FL to offset the bias introduced by non-IID data and accelerate the convergence speed. Xu et al. [141] proposes a stochastic optimization problem that jointly connects client selection and bandwidth allocation under long-term client constraints. Lai et al. [142] proposed Oort to improve the performance of joint training and testing by guiding participant selection. Oort uses existing data to improve model accuracy and speed up training. Wu et al. [143] proposed a probabilistic node selection framework, FedPNS, which dynamically changes the probability of each node being selected based on the output of the optimal aggregation. Cho et al. [144] presents a convergence analysis of FL of biased customer selection via efficient selection aggregation and quantifies how bias affects the convergence speed. Li et al. [145] proposed PyramidFL to speed up FL training while achieving higher final model performance. At the heart of PyramidFL is fine-grained client selection, not only focusing on the differences in client selection between those selected and unselected participants, but also fully exploiting data and system heterogeneity within selected clients. Ribero et al. [46] proposes a federated averaging algorithm assisted by an adaptive sampling technique, F3AST, an unbiased algorithm that dynamically learns an availability-dependent client selection strategy that asymptotically minimizes client sampling variance to global model convergence The impact improves the performance of joint learning.
@@ III-C2 Clustered Federated Learning
Clustering FL (CFL) is used to deal with situations where the data distribution is highly heterogeneous. This approach trains the model in a more homogeneous group by clustering clients based on the similarity of their data. CFL [36] is a method of recursively clustering clients by exploiting the cosine similarity between client gradient updates during each round of FL training. This clustering allows for a more targeted and efficient distribution of training tasks among clients with similar characteristics.
IFCA [146] is another clustered FL method that utilizes the loss of models within different clusters on the client’s local dataset to cluster clients. In each round, IFCA randomly selects a portion of clients within each cluster for training and updates the intra-cluster models by aggregating the local models within the cluster. PACFL [147] clusters clients by uploading data information from each client to the server and taking measurements of the principal angles between client data subspaces before the FL process begins. During each round of joint learning, PACFL updates each intra-cluster model with a weighted aggregation of the local models of the intra-cluster clients, enabling personalized learning while still taking advantage of clustering. FedSoft [148] designed to train both local personalized models and high-quality cluster models. To limit the workload of the clients, FedSoft employs proximal updating, which allows only a subset of clients to complete an optimization task in each round of communication. ACFL [35] introduces a mean-shift clustering algorithm that intelligently groups clients based on their local data distributions. This clustering approach allows FL systems to identify groups of clients with similar data characteristics, enabling more targeted and efficient model training.
III-C3 Hierarchical Federated Learning